OpenStudy (anonymous):

Solve |2x + 8| = 6. Show its equivalent compound sentence and show its solution

OpenStudy (anonymous):

| 2x + 8 | = 6 means: 2x + 8 = 6 or 2x + 8 = -6 2x = -2 or.....2x = -14 x = -1 or x = -7 two solutions: x = -1 or x = -7

OpenStudy (anonymous):

thank you